Add.
(r-4s+8)+(8r+3s)+(s+7)

Answer by lenny460(1073)

(r - 4s + 8) + (8r + 3s) + (s + 7)
r - 4s + 8 + 8r + 3s + s + 7
Combine like terms:
r + 8r = 9r
-4s + 3s = -1s
-1s + s = 0
8 + 7 = 15
The Answer:
9r + 15